OpenStudy (anonymous):

which method is used to solve h^2 + 4h + 7 = 0

hero (hero):

@leonela0131040, have you tried solving it? If so, which method would you use to solve it?

OpenStudy (anonymous):

\[h^2+4h+7=0\\\Delta=4^2-4.1.7\\\Delta=-12\\\\h=\frac{-4\pm\sqrt{-12}}{2}\\h=\frac{-4\pm2i\sqrt3}{2}\\\\h'=-2+i\sqrt3\\h''=-2-i\sqrt3\]